OpenStudy (steven513):

The equation of the motion of a body suspended by a spring is T=2π√m/k, where T is the time period, m is the mass of the body, and k is the force constant of spring. use this information to find the time period if a body of mass 10 kg is suspended by a spring of force constant 1400 Nm^-1. Round to the nearest tenth. My answer: I did not exactly complete the problem as I am unsure if these values belong there; T=2π√10/1400, is this the correct step or was I supposed to do something else? Thank You

OpenStudy (johnweldon1993):

Nope that was all you had to do \[\large T = 2\pi \sqrt{\frac{m}{k}}\] Mass is 10 kg Spring Constant is 1400 N/m Therefor: \[\large T = 2\pi \sqrt{\frac{10}{1400}}\] *Always make sure the units are correct! \[\large rads\times \sqrt{\frac{\cancel{kg}}{\frac{\cancel{kg} \cdot \cancel{m}}{s^2 \cdot \cancel{m}}}} = \frac{rads}{s}\] Which is a correct unit that we need

OpenStudy (steven513):

Thank You @johnweldon1993

OpenStudy (steven513):

would there be a need to simply that equation? @johnweldon1993